### v1

@@ 2. $(ii)$
[EGA4_2, Proposition 6.11.6] gives $(i)$. For $(ii)$, it suffices to show for a finite type inclusion of Noetherian domains $R\subseteq B$, if $R$ is $(S_{2})$, then $B_{g}$ is $(S_{2})$ for some $g\in B^{\circ}$. To see this, we may assume $R\to B$ is flat [stacks, Tag 051R], and has Cohen–Macaulay fibers [stacks, Tag 045U]. Then $B$ is $(S_{2})$ by [stacks, Tag 0339]. ∎
@@
    For every finite ring map $R\to B$ where $B$ is an integral domain, there exists a finite sub-$B$-algebra $C$ of $B^{\sigma}$ that is $(S_{2})$.
@@ 3. $(iii)$
As $(ii)$ implies $(i)$, it suffices to show $(iii)$ implies $(ii)$. Assume $(iii)$ and assume $R$ is an integral domain. We must show that there exists a finite sub-$R$-algebra $R^{\prime}$ of $R^{\sigma}$ that is $(S_{2})$. Let $U$ be the $(S_{2})$ locus of $R$, which is open by Lemma 7.12, and let $\mathfrak{p}\not\in U$, so $\operatorname{ht}(\mathfrak{p})>1$.
The local ring $R_{\mathfrak{p}}$ is semi-Nagata by Theorem 7.8, so by Lemma 4.2 there exists a finite sub-$R$-algebra $R_{1}$ of $R^{\sigma}$ such that $(R_{1})_{\mathfrak{p}}$ is $(S_{2})$. As $R$ is $(S_{2})$-2, the $(S_{2})$ locus of the ring $R_{1}$ is open by Lemma 7.12. As the constructible topology of $\operatorname{Spec}(R_{1})$ is compact [stacks, Tag 0901], there exists $f_{1}\in R\setminus\mathfrak{p}$ such that $(R_{1})_{f_{1}}$ is $(S_{2})$. Since $R_{1}\subseteq R^{\sigma}$ we know $R=R_{1}$ over $U$, so the image $Z_{1}$ of the non-$(S_{2})$ locus of $R_{1}$ in $\operatorname{Spec}(R)$ is disjoint from $U\cup D(f_{1})$. If $Z_{1}\neq\emptyset$, take $\mathfrak{q}\in Z_{1}$, then similarly the semilocal ring $(R_{1})_{\mathfrak{q}}$ is semi-Nagata, and we can find $R_{2}\subseteq R_{1}^{\sigma}\subseteq R^{\sigma}$ (Lemma 4.3) so that the image $Z_{2}$ of the non-$(S_{2})$ locus of $R_{2}$ in $\operatorname{Spec}(R)$ is disjoint from $U\cup D(f_{1})\cup D(f_{2})$ and $f_{2}\in R\setminus\mathfrak{q}$. As the topological space $\operatorname{Spec}(R)$ is Noetherian, we get our desired $R^{\prime}$ after finitely many steps. ∎
@@
Let $R$ be a Noetherian integral domain that satisfies $(iii)$. We want to show there exists a finite sub-$R$-algebra of $R^{\sigma}$ that is $(S_{2})$.
Let $\mathfrak{m}\in\operatorname{Max}(R)$. Then $R_{\mathfrak{m}}$ is semi-Nagata by Theorem 7.8, so by Lemma 4.2 there exists a finite sub-$R$-algebra $R(\mathfrak{m})$ of $R^{\sigma}$ such that $(R(\mathfrak{m}))_{\mathfrak{m}}$ is $(S_{2})$.
@@
Take finitely many $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}$ so that $D(f(\mathfrak{m}_{i}))\ (1\leq i\leq n)$ cover $\operatorname{Spec}(R)$ and let $R_{1}$ be the $R$-algebra generated by all $R(\mathfrak{m}_{i})$. Then $R(\mathfrak{m}_{i})_{f(\mathfrak{m}_{i})}\subseteq(R_{1})_{f(\mathfrak{m}_{i})}$, so by Remark 4.7 and Lemma 6.1 we have $\operatorname{OS}_{2}^{\wedge}({R_{1}})=\emptyset$. As $R_{1}^{\sigma}\subseteq R^{\sigma}$ (Lemma 4.3) we may replace $R$ by $R_{1}$ to assume $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$. Apply the same construction again, we see from Theorem 4.8$(iii)$ (and Lemma 4.2) that $R(\mathfrak{m}_{i})_{f(\mathfrak{m}_{i})}=R^{\sigma}_{f(\mathfrak{m}_{i})}$ for all $i$, so $R^{\sigma}$ is finite over $R$. ∎
@@
[EGA4_2, Proposition 6.3.8] tells us a Cohen–Macaulay ring is an $(S_{1})$-ring, and [EGA4_2, Proposition 6.11.8 and Remarques 6.11.9] tell us a Cohen–Macaulay ring is $(S_{2})$-2. ∎
@@
Let $R$ be a semi-Nagata ring that is $(S_{1})$ and has irreducible spectrum. Let $F$ be the total fraction ring of $R$ and let $M$ be a finite sub-$R$-module of $F$. Assume $\operatorname{OS}_{2}^{\wedge}({R_{\operatorname{red}}})=\emptyset$. Then $M^{n\sigma}:=\bigcap_{\mathfrak{p}\in\operatorname{Spec}_{1}(R)}M_{\mathfrak{p}}$ is finite and $(S_{2})$.
@@
Let $U$ be the locus where $M$ is $(S_{2})$. Then $U$ is open (Lemma 7.12), and contains $\operatorname{Spec}_{1}(R)$ as $R$ is $(S_{1})$. Let $j:U\to\operatorname{Spec}(R)$ be the canonical open immersion. It is clear that $M^{n\sigma}=j_{*}j^{*}M$.
@@
Let $\mathfrak{p}_{0}$ be the minimal prime of $R$, and let $\overline{R}=R/\mathfrak{p}_{0}=R_{\operatorname{red}}$. Note that $\operatorname{Ass}_{R}(M)=\{\mathfrak{p}_{0}\}$. For $\mathfrak{p}\in\operatorname{Spec}(R)\setminus U$, the completion $\overline{R}^{\wedge}_{\mathfrak{p}}$ is $(S_{1})$ as $\overline{R}$ is $(S_{1})$ with $(S_{1})$ formal fibers. Moreover, for every $P_{0}\in\operatorname{Min}(\overline{R}^{\wedge}_{\mathfrak{p}})$, we have $\dim(\overline{R}^{\wedge}_{\mathfrak{p}}/P_{0})>1$ as $\operatorname{ht}(\mathfrak{p})>1$ and as $\operatorname{OS}_{2}^{\wedge}({\overline{R}_{\mathfrak{p}}})=\emptyset$ (Corollary 6.5). Therefore for all $P_{0}\in\operatorname{Ass}(\overline{R}^{\wedge}_{\mathfrak{p}})$, we have $\dim(\overline{R}^{\wedge}_{\mathfrak{p}}/P_{0})>1$. By [stacks, Tag 0BK3], $M^{n\sigma}=j_{*}j^{*}M$ is finite. It is $(S_{2})$ by for example [EGA4_2, Théorème 5.10.5]. ∎
@@ 8. The local lifting argument
In this section, we present an adapted version of Nishimura’s argument for local lifting [Nishimura-semilocal-lifting]. It is a variant of Rotthaus’ argument [Rotthaus-qe-semilocal-lifting], which is axiomitized in [BI-semilocal-lifting]. A key component of the argument in all the three aforementioned articles is that the property of concern must imply reducedness (cf. [BI-semilocal-lifting, Remark after Theorem 2.3]). We remove this restriction.
Discussion 8.1.
@@
A Noetherian ring $A$ is said to be a $\mathbf{P}$-ring if its formal fibers are geometrically $\mathbf{P}$, in other words, $A_{\mathfrak{p}}\to A^{\wedge}_{\mathfrak{p}}$ is a $\mathbf{P}$-map for all $\mathfrak{p}\in\operatorname{Spec}(A)$. By [stacks, Tag 0BIU], if $\mathbf{P}$ satisfies (I)(III)(IV) below, then a Noetherian ring $A$ is a $\mathbf{P}$-ring if and only if $A_{\mathfrak{m}}\to A^{\wedge}_{\mathfrak{m}}$ is a $\mathbf{P}$-map for all $\mathfrak{m}\in\operatorname{Max}(A)$. When $A$ is semilocal, this is to say $A\to A^{\wedge}$ is a $\mathbf{P}$-map (cf. [EGA4_2, Proposition 7.3.14]); and if $\mathbf{P}$ satisfies (I)(II)(III)(IV) below, this is also equivalent to that for every finite $A$-algebra $B$ that is an integral domain, $(B^{\circ})^{-1}B^{\wedge}$ satisfies $\mathbf{P}$, see [EGA4_2, Proposition 7.3.16]. By [stacks, Tag 0BIV], an essentially finitely generated algebra over a $\mathbf{P}$-ring is a $\mathbf{P}$-ring.
@@
Whether or not $\mathbf{P}$ satisfies (VI) is generally called the Grothendieck localization problem for $\mathbf{P}$. The paper [Mur-Grothendieck-localization] provides a uniform treatment to this problem, and provides a list of references of known results on what properties satisfy (I)(VI). In particular, (I)(VI) hold for $\mathbf{P}$=“$(S_{k})$,” “Cohen–Macaulay,” “Gorenstein,” and “lci.”
@@
In [Nishimura-semilocal-lifting, BI-semilocal-lifting], $\mathbf{Q}$=“reduced,” and $\mathfrak{c}(A)$ is the unique radical ideal that satisfies $V(\mathfrak{c}(A))=\operatorname{Spec}(A)\setminus U_{\mathbf{P}}(A)$. The lifting of $\mathbf{Q}$-rings is [Marot-Nagata-Lift]. As a reduced ring is $(R_{0})$ we see $\mathfrak{c}(A)_{\mathfrak{q}}=A_{\mathfrak{q}}$ for all $\mathfrak{q}\in\operatorname{Min}(A)$, in particular $\mathfrak{c}(A)\neq 0$. We have $\varphi(\mathfrak{c}(A))B=\mathfrak{c}(B)$ for all $\varphi:A\to B$ in $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}-1}$ as $\operatorname{Spec}(\varphi)^{-1}(U_{\mathbf{P}}(A))=U_{\mathbf{P}}(B)$ and as the fibers of $\varphi$ are reduced, so $\varphi(\mathfrak{c}(A))B$ is radical.
In our case, we do not have such luxury, and it is necessary to find the assignments case-by-case. We will find assignments on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ for $\mathbf{Q}$ trivial and $\mathbf{P}$=“$(S_{1})$,” $\mathbf{Q}$=“$(S_{1})$” and $\mathbf{P}$=“$(S_{2})$,” and $\mathbf{Q}$=“‘$(S_{2})$” and $\mathbf{P}$=“$(S_{k})$$(k\geq 3)$, “Gorenstein,” and “lci.”
@@
Write $C=\mathfrak{c}(R^{\wedge})\neq 0$, and for every $n\in\mathbf{Z}_{\geq 1}$, $C_{n}=C+I^{n}R^{\wedge}$, $\mathfrak{a}_{n}=C_{n}\cap R$. We will show $C_{n}=\mathfrak{a}_{n}R^{\wedge}$ for all $n$, which implies $C\cap R\neq 0$ by [Rotthaus-qe-semilocal-lifting, Lemma 2], which then tells us the generic fiber $(R^{\circ})^{-1}R^{\wedge}$ is $\mathbf{P}$, as $V(C)=\operatorname{Spec}(R^{\wedge})\setminus U_{\mathbf{P}}(R^{\wedge})$.
@@
We know $(R^{\wedge})_{\mathfrak{p}}$ is (quasi-)excellent as $R^{\wedge}$ is complete, therefore $((R^{\wedge})_{\mathfrak{p}})^{*}$ is quasi-excellent [formal-lifting-excellence-Gabber]. In particular, both $(R^{\wedge})_{\mathfrak{p}}$ and $((R^{\wedge})_{\mathfrak{p}})^{*}$ are $\mathbf{P}$-rings. By (8) $(R_{\mathfrak{p}})^{*}$ is also.

### v2

@@ 2. $(ii)$
[EGA ̵IV2, Proposition 6.11.6] gives $(i)$. For $(ii)$, it suffices to show for a finite type inclusion of Noetherian domains $R\subseteq B$, if $R$ is $(S_{2})$, then $B_{g}$ is $(S_{2})$ for some $g\in B^{\circ}$. To see this, we may assume $R\to B$ is flat [Stacks, Tag 051R], and has Cohen–Macaulay fibers [Stacks, Tag 045U]. Then $B$ is $(S_{2})$ by [Stacks, Tag 0339]. ∎
@@
    For every finite ring map $R\to B$ where $B$ is an integral domain, there exists a finite subalgebra $C$ of $B^{\sigma}$ that is $(S_{2})$.
@@ 3. $(iii)$
As $(ii)$ implies $(i)$, it suffices to show $(iii)$ implies $(ii)$. Assume $(iii)$ and assume $R$ is an integral domain. We must show that there exists a finite subalgebra $R^{\prime}$ of $R^{\sigma}$ that is $(S_{2})$. Let $U$ be the $(S_{2})$ locus of $R$, which is open by Lemma 7.13, and let $\mathfrak{p}\not\in U$, so $\operatorname{ht}(\mathfrak{p})>1$.
The local ring $R_{\mathfrak{p}}$ is semi-Nagata by Theorem 7.8, so by Lemma 4.2 there exists a finite subalgebra $R_{1}$ of $R^{\sigma}$ such that $(R_{1})_{\mathfrak{p}}$ is $(S_{2})$. As $R$ is $(S_{2})$-2, the $(S_{2})$ locus $U_{1}$ of the ring $R_{1}$ is open by Lemma 7.13, and we have $\bigcap_{h\in R\setminus\mathfrak{p}}D_{R_{1}}(f)\subseteq U_{1}$. As the constructible topology of $\operatorname{Spec}(R_{1})$ is compact [Stacks, Tag 0901], we see there exists $f_{1}\in R\setminus\mathfrak{p}$ such that $(R_{1})_{f_{1}}$ is $(S_{2})$. Since $R_{1}\subseteq R^{\sigma}$ we know $R=R_{1}$ over $U$, so the image $Z_{1}$ of the non-$(S_{2})$ locus of $R_{1}$ in $\operatorname{Spec}(R)$ is disjoint from $U\cup D(f_{1})$. If $Z_{1}\neq\emptyset$, take $\mathfrak{q}\in Z_{1}$, then similarly the semilocal ring $(R_{1})_{\mathfrak{q}}$ is semi-Nagata, and we can find $R_{2}\subseteq R_{1}^{\sigma}\subseteq R^{\sigma}$ (Lemma 4.3) so that the image $Z_{2}$ of the non-$(S_{2})$ locus of $R_{2}$ in $\operatorname{Spec}(R)$ is disjoint from $U\cup D(f_{1})\cup D(f_{2})$ and $f_{2}\in R\setminus\mathfrak{q}$. As the topological space $\operatorname{Spec}(R)$ is Noetherian, we get our desired $R^{\prime}$ after finitely many steps. ∎
@@
Let $R$ be a Noetherian integral domain that satisfies $(iii)$. We want to show there exists a finite subalgebra of $R^{\sigma}$ that is $(S_{2})$.
Let $\mathfrak{m}\in\operatorname{Max}(R)$. Then $R_{\mathfrak{m}}$ is semi-Nagata by Theorem 7.8, so by Lemma 4.2 there exists a finite subalgebra $R(\mathfrak{m})$ of $R^{\sigma}$ such that $(R(\mathfrak{m}))_{\mathfrak{m}}$ is $(S_{2})$.
@@
Take finitely many $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}$ so that $D(f(\mathfrak{m}_{i}))\ (1\leq i\leq n)$ cover $\operatorname{Spec}(R)$ and let $R_{1}$ be the $R$-algebra generated by all $R(\mathfrak{m}_{i})$. Then $R(\mathfrak{m}_{i})_{f(\mathfrak{m}_{i})}\subseteq(R_{1})_{f(\mathfrak{m}_{i})}$, so by Remark 4.7 and Lemma 6.1 we have $\operatorname{OS}_{2}^{\wedge}({R_{1}})=\emptyset$. As $R_{1}^{\sigma}\subseteq R^{\sigma}$ (Lemma 4.3) we may replace $R$ by $R_{1}$ to assume $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$. Apply the same construction again, we see from Theorem 4.8(iii)(iii) (and Lemma 4.2) that $R(\mathfrak{m}_{i})_{f(\mathfrak{m}_{i})}=R^{\sigma}_{f(\mathfrak{m}_{i})}$ for all $i$, so $R^{\sigma}$ is finite over $R$. ∎
@@
[EGA ̵IV2, Proposition 6.3.8] tells us a Cohen–Macaulay ring is an $(S_{1})$-ring, and [EGA ̵IV2, Proposition 6.11.8 and Remarques 6.11.9] tell us a Cohen–Macaulay ring is $(S_{2})$-2. ∎
@@
Let $R$ be a semi-Nagata ring. Assume $\operatorname{OS}_{2}^{\wedge}({R/\mathfrak{p}_{0}})=\emptyset$ for all $\mathfrak{p}_{0}\in\operatorname{Min}(R)$.
Let $\Sigma=\{\mathfrak{p}\in\operatorname{Spec}(R)\mid\operatorname{ht}(\mathfrak{p}/\mathfrak{p}_{0})=1\text{ for some }\mathfrak{p}_{0}\in\operatorname{Min}(R)\}$. Assume that $\Sigma=\operatorname{Spec}_{1}(R)$.
Let $M$ be a finite module so that $\operatorname{Ass}(M)=\operatorname{Min}(R)$. Then $N:=\bigcap_{\mathfrak{p}\in\Sigma}M_{\mathfrak{p}}$ is finite and $(S_{2})$.
Note that the assumptions on $\Sigma$ and $M$ are satisfied, for example, when $R$ is an integral domain and $M$ is torsion-free.
@@
Let $U$ be the locus where $M$ is $(S_{2})$. Then $U$ is open (Lemma 7.13), and contains $\Sigma$ as $\Sigma=\operatorname{Spec}_{1}(R)$ and $\operatorname{Ass}(M)=\operatorname{Min}(R)$. Let $j:U\to\operatorname{Spec}(R)$ be the canonical open immersion. It is now clear that $N=j_{*}j^{*}M$.
Let $\mathfrak{p}_{0}$ be a minimal prime of $R$, and let $\overline{R}=R/\mathfrak{p}_{0}$. By [Stacks, Tag 0BK3] to show $N$ is finite it suffices to show for $\mathfrak{p}\in\operatorname{Spec}(R)\setminus U$ containing $\mathfrak{p}_{0}$, every $P_{0}\in\operatorname{Ass}(\overline{R}^{\wedge}_{\mathfrak{p}})$ satisfies $\dim(\overline{R}^{\wedge}_{\mathfrak{p}}/P_{0})>1$.
@@
We know $\overline{R}^{\wedge}_{\mathfrak{p}}$ is $(S_{1})$ as $\overline{R}$ is $(S_{1})$ with $(S_{1})$ formal fibers. Moreover, for every $P_{0}\in\operatorname{Min}(\overline{R}^{\wedge}_{\mathfrak{p}})$, we have $\dim(\overline{R}^{\wedge}_{\mathfrak{p}}/P_{0})>1$ as $\operatorname{ht}(\mathfrak{p})>1$ and as $\operatorname{OS}_{2}^{\wedge}({\overline{R}_{\mathfrak{p}}})=\emptyset$ (Corollary 6.5). Thus for all $P_{0}\in\operatorname{Ass}(\overline{R}^{\wedge}_{\mathfrak{p}})$, we have $\dim(\overline{R}^{\wedge}_{\mathfrak{p}}/P_{0})>1$. Therefore $N$ is finite. It is $(S_{2})$ by for example [EGA ̵IV2, Théorème 5.10.5]. ∎
8. Lifting the semi-Nagata property
In this section, we prove the following result, giving a partial answer to Question 1.1 for the semi-Nagata property.
Theorem 8.1.
Let $R$ be a Noetherian ring, $I$ an ideal of $R$. Assume that
  1. (1)
    $R$ is $I$-adically complete.
  2. (2)
    $R/I$ is semi-Nagata.
  3. (3)
    $R$ is universally catenary.
Then $R$ is semi-Nagata.
We proceed with the proof. We use the characterization Theorem 7.14 without further mentioning. By Noetherian induction, we may assume
  1. (4)
    $R/\mathfrak{a}$ is semi-Nagata for all nonzero ideals $\mathfrak{a}$ of $R$.
By Definition 7.1 we may also assume
  1. (5)
    $R$ is an integral domain,
and we only need to find a finite subalgebra $R^{\prime}$ of $K$, the fraction field of $R$, that is $(S_{2}).$ As $R$ is $I^{\prime}$-adically complete for all ideals $I^{\prime}\subseteq I$, by (4) we may assume
  1. (6)
    II is generated by a single element $f\neq 0$.
Let $\mathcal{M}$ be the set of minimal prime divisors of $I$.
We will construct a sequence of submodules $R=M_{0}\subseteq M_{1}\subseteq\ldots$ of $K$ so that the union $M=\bigcup_{i}M_{i}$ is finite over $R$ and that $M/fM$ is $(S_{1})$. We will later show that the existence of such an $M$ implies $R$ is semi-Nagata.
For a finite submodule $X$ of $K$ we let $J(X)$ be the intersection of the embedded primes of the $R$-module $X/fX$. The module $H^{0}_{J(X)}(X/fX)$ is canonically identified with $H(X):=H^{1}_{J(X)}(X)[f]$ as $f$ is a nonzerodivisor on $X$, and we have a module $X^{+}:=X+_{J(X)}H(X)\subseteq K$ fitting into an exact sequence
[…]
similar to the construction $R+_{I}M$ in Discussion 5.1. We will show that $M_{0}=R$ and $M_{i+1}=M_{i}^{+}$ gives the desired sequence of modules.
The exact sequence above gives an exact sequence
$X+/fX+{X^{+}/fX^{+}}H(X){H(X)}0.{0.}$
As $H(X)$ is identified with $H^{0}_{J(X)}(X/fX)$ we see from primary decomposition that $\operatorname{Ass}\left(\frac{X/fX}{H(X)}\right)$ is the set of minimal prime divisors of $X/fX$. We know $\operatorname{Supp}(X)=\operatorname{Spec}(R)$ as $X$ is a submodule of $K$, so $\operatorname{Supp}(X/fX)=\operatorname{Spec}(R/fR)$ by Nakayama’s Lemma, thus $\operatorname{Ass}\left(\frac{X/fX}{H(X)}\right)=\mathcal{M}$. It follows that the associated primes of the module $M/fM=\operatorname{colim}_{i}M_{i}/fM_{i}=\operatorname{colim}_{i}\frac{M_{i}/fM_{i}}{H(M_{i})}$ are all in $\mathcal{M}$, so $M/fM$ is $(S_{1})$ as soon as $M/fM$ is finite.
The discussion above also shows $\operatorname{Ass}(X^{+}/fX^{+})\subseteq\mathcal{M}\cup V(J(X))$, so $J(X^{+})\supseteq J(X)$. As $R$ is Noetherian, there exists $i_{0}$ so that $J(M_{i})=J(M_{i_{0}})$ for all $i\geq i_{0}$. Let $\mathcal{J}$ be the set of minimal prime divisors of $J:=J(M_{i_{0}})$. For every $i\geq i_{0}$ we have an exact sequence
[…]
which gives an injection $H(M_{i+1})=H^{0}_{J}(M_{i+1}/fM_{i+1})\to H(M_{i})$ as $H^{0}_{J}\left(\frac{M_{i}/fM_{i}}{H(M_{i})}\right)=0$. Thus there exists $i_{1}\geq i_{0}$ so that $H(M_{i+1})_{\mathfrak{P}}=H(M_{i})_{\mathfrak{P}}$ for all $\mathfrak{P}\in\mathcal{J}$ and all $i\geq i_{1}$, as the lengths of $H(M_{i})_{\mathfrak{P}}$ are finite. It follows that
$\left(\frac{M_{i_{1}}/fM_{i_{1}}}{H(M_{i_{1}})}\right)_{\mathfrak{P}}=\left(\frac{M_{i_{1}+1}/fM_{i_{1}+1}}{H(M_{i_{1}+1})}\right)_{\mathfrak{P}}=\ldots=(M/fM)_{\mathfrak{P}}$
for all $\mathfrak{P}\in\mathcal{J}$; note that the same is true for all $\mathfrak{P}\in D(J)$, as, in that case, $(M_{i})_{\mathfrak{P}}=(M_{i+1})_{\mathfrak{P}}$.
We now apply Theorem 7.17 to the semi-Nagata ring $A=R/fR$ and the module $Y=\frac{M_{i_{1}}/fM_{i_{1}}}{H(M_{i_{1}})}$. The condition $\operatorname{Ass}(Y)=\operatorname{Min}(A)$ follows from the construction. The conditions $\operatorname{OS}_{2}^{\wedge}({A/\mathfrak{p}_{0}})=\emptyset$ and $\Sigma=\operatorname{Spec}_{1}(A)$ follows from the condition $R$ is universally catenary, see Remark 4.6. If $\mathfrak{P}\in D(J)$ or $\mathfrak{P}\in\mathcal{J}$ then we know $(M/fM)_{\mathfrak{P}}=Y_{\mathfrak{P}}$. However, $D(J)\cup\mathcal{J}$ covers $\operatorname{Spec}_{1}(A)$ as $R$ is catenary. We see $(M/fM)_{\mathfrak{P}}=Y_{\mathfrak{P}}$ for all $\mathfrak{P}\in\operatorname{Spec}_{1}(A)$, so Theorem 7.17 tells us $M/fM$ is finite, hence $(S_{1})$ as noted before.
As $M$ is a submodule of $K$, we know $M\subseteq M_{\mathfrak{p}}$ for all $\mathfrak{p}\in\mathcal{M}$, in particular $M$ is $f$-adically separated. Therefore $M$ is a submodule of its $f$-adic completion, which is finite as $R$ is $f$-adically complete and $M/fM$ is finite.
We have found a finite submodule $M$ of $K$ containing $R$ so that $M/fM$ is $(S_{1})$. Let $\mathfrak{p}\in V(f)$. The fibers of $R_{\mathfrak{p}}/fR_{\mathfrak{p}}\to R_{\mathfrak{p}}^{\wedge}/fR_{\mathfrak{p}}^{\wedge}$ are $(S_{1})$ as $R/fR$ is semi-Nagata, so $M_{\mathfrak{p}}^{\wedge}/fM_{\mathfrak{p}}^{\wedge}$ is $(S_{1})$ by [EGA ̵IV2, Proposition 6.4.1], thus $M_{\mathfrak{p}}^{\wedge}$ is $(S_{1})$ by [EGA ̵IV2, Proposition 3.4.4]. We know $R_{g}=M_{g}$ for some $g\in R^{\circ}$, so $(R_{\mathfrak{p}}^{\wedge})_{g}=(M_{\mathfrak{p}}^{\wedge})_{g}$ and $g$ is a nonzerodivisor in both $R_{\mathfrak{p}}^{\wedge}$ and $M_{\mathfrak{p}}^{\wedge}$, showing that $R_{\mathfrak{p}}^{\wedge}$ is $(S_{1})$.
As $\operatorname{OS}_{2}^{\wedge}({R_{\mathfrak{p}}})=\emptyset$ (Remark 4.6), by Theorem 4.8 there exists a finite subalgebra $R(\mathfrak{p})$ of $R^{\sigma}$ so that $R(\mathfrak{p})_{\mathfrak{p}}$ is $(S_{2})$. Note that this is the same as $R(\mathfrak{p})_{\mathfrak{p}}$ is an $(S_{2})$ $R_{\mathfrak{p}}$-module, as $R$ is universally catenary (Remark 7.11). As $R/\mathfrak{p}$ is semi-Nagata and therefore $(S_{2})$-2, the proof of [EGA ̵IV2, Proposition 6.11.6] tells us there exists $h(\mathfrak{p})\not\in\mathfrak{p}$ so that $R(\mathfrak{p})_{\mathfrak{P}}$ is $(S_{2})$ for all $\mathfrak{P}\in V(\mathfrak{p})\cap D(h)$. As the constructible topology of $R/fR$ is compact [Stacks, Tag 0901] we see there exist finitely many finite subalgebras $R_{1},\ldots,R_{n}$ of $R^{\sigma}$ so that for every $\mathfrak{P}\in V(f)$ there exists a $j$ such that $(R_{j})_{\mathfrak{P}}$ is $(S_{2})$, in other words, $(R_{j})_{\mathfrak{P}}=(R^{\sigma})_{\mathfrak{P}}$ (Theorem 4.8). Let $R^{\prime}$ be the finite subalgebra generated by all $R_{j}$, so $R^{\prime}_{\mathfrak{P}}=(R^{\sigma})_{\mathfrak{P}}$ for all $\mathfrak{P}\in V(f)$, thus $R^{\prime}_{\mathfrak{P}}$ is $(S_{2})$ for all $\mathfrak{P}\in V(f)$. As $R$ is $f$-adically complete, this tells us $R^{\prime}_{\mathfrak{P}}$ is $(S_{2})$ for all $\mathfrak{P}\in\operatorname{Max}(R)$, so $R^{\prime}$ is $(S_{2})$, as desired.
@@ 9. The local lifting argument
In this section, we present an adapted version of Nishimura’s argument for local lifting [Nis81]. It is a variant of Rotthaus’ argument [Rot79], which is axiomitized in [BI84]. A key component of the argument in all the three aforementioned articles is that the property of concern must imply reducedness (cf. [BI84, Remark after Theorem 2.3]). We remove this restriction.
Discussion 9.1.
@@
A Noetherian ring $A$ is said to be a $\mathbf{P}$-ring if its formal fibers are geometrically $\mathbf{P}$, in other words, $A_{\mathfrak{p}}\to A^{\wedge}_{\mathfrak{p}}$ is a $\mathbf{P}$-map for all $\mathfrak{p}\in\operatorname{Spec}(A)$. By [Stacks, Tag 0BIU], if $\mathbf{P}$ satisfies (I)(III)(IV) below, then a Noetherian ring $A$ is a $\mathbf{P}$-ring if and only if $A_{\mathfrak{m}}\to A^{\wedge}_{\mathfrak{m}}$ is a $\mathbf{P}$-map for all $\mathfrak{m}\in\operatorname{Max}(A)$. When $A$ is semilocal, this is to say $A\to A^{\wedge}$ is a $\mathbf{P}$-map (cf. [EGA ̵IV2, Proposition 7.3.14]); and if $\mathbf{P}$ satisfies (I)(II)(III)(IV) below, this is also equivalent to that for every finite $A$-algebra $B$ that is an integral domain, $(B^{\circ})^{-1}B^{\wedge}$ satisfies $\mathbf{P}$, see [EGA ̵IV2, Proposition 7.3.16]. By [Stacks, Tag 0BIV], an essentially finitely generated algebra over a $\mathbf{P}$-ring is a $\mathbf{P}$-ring.
@@
Whether or not $\mathbf{P}$ satisfies (VI) is generally called the Grothendieck localization problem for $\mathbf{P}$. The paper [Mur22] provides a uniform treatment to this problem, and provides a list of references of known results on what properties satisfy (I)(VI). In particular, (I)(VI) hold for $\mathbf{P}$$(S_{k})$,” “Cohen–Macaulay,” “Gorenstein,” and “lci.”
@@
In [Nis81, BI84], $\mathbf{Q}$=“reduced,” and $\mathfrak{c}(A)$ is the unique radical ideal that satisfies $V(\mathfrak{c}(A))=\operatorname{Spec}(A)\setminus U_{\mathbf{P}}(A)$. The lifting of $\mathbf{Q}$-rings is [Mar75]. As a reduced ring is $(R_{0})$ we see $\mathfrak{c}(A)_{\mathfrak{q}}=A_{\mathfrak{q}}$ for all $\mathfrak{q}\in\operatorname{Min}(A)$, in particular $\mathfrak{c}(A)\neq 0$. We have $\varphi(\mathfrak{c}(A))B=\mathfrak{c}(B)$ for all $\varphi:A\to B$ in $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}-1}$ as $\operatorname{Spec}(\varphi)^{-1}(U_{\mathbf{P}}(A))=U_{\mathbf{P}}(B)$ and as the fibers of $\varphi$ are reduced, so $\varphi(\mathfrak{c}(A))B$ is radical.
In our case, we do not have such luxury, and it is necessary to find the assignments case-by-case. We will find assignments on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ for $\mathbf{Q}$ trivial and $\mathbf{P}$$(S_{1})$,” $\mathbf{Q}$$(S_{1})$” and $\mathbf{P}$$(S_{2})$,” and $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$$(S_{k})$$(k\geq 3)$, “Gorenstein,” and “lci.”
@@
Write $C=\mathfrak{c}(R^{\wedge})\neq 0$, and for every $n\in\mathbf{Z}_{\geq 1}$, $C_{n}=C+I^{n}R^{\wedge}$, $\mathfrak{a}_{n}=C_{n}\cap R$. We will show $C_{n}=\mathfrak{a}_{n}R^{\wedge}$ for all $n$, which implies $C\cap R\neq 0$ by [Rot79, Lemma 2], which then tells us the generic fiber $(R^{\circ})^{-1}R^{\wedge}$ is $\mathbf{P}$, as $V(C)=\operatorname{Spec}(R^{\wedge})\setminus U_{\mathbf{P}}(R^{\wedge})$.
@@
We know $(R^{\wedge})_{\mathfrak{p}}$ is (quasi-)excellent as $R^{\wedge}$ is complete, therefore $((R^{\wedge})_{\mathfrak{p}})^{*}$ is quasi-excellent [KS21]. In particular, both $(R^{\wedge})_{\mathfrak{p}}$ and $((R^{\wedge})_{\mathfrak{p}})^{*}$ are $\mathbf{P}$-rings. By (8) $(R_{\mathfrak{p}})^{*}$ is also.